Grade 11Inorganic Chemistry

canyou pleasegivve answer for this”pure o2 diffuses through an aperture in 224 seconds, wheres mixture of o2 and anther gas containing 80% o2 diffuses from the same in 234sec what is the molecular weight of gas?”please mail the answer to the given mail.

shaik mohaseen
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
I can't mail you as you have asked here.
 

For gaseous mixture 80% O2, 20% of other gas m   

Average molecular weight Mm = 32 (for O2) × 80 + 20 × m / 100   

Now, for diffusion of gaseous mixture and pure O2   

rO2 / rm = √Mm / MO2 or VO2 / EO2 × Em / Vm = √Mm / MO2   

∴ 1 / 224 × 234 / 1 = √Mm / 32 

∴ Mm= 34.92  

∴ 32 × 80 + 20 × m / 100 = 34.92   

∴ m = 46.6